# Improving Expressivity of Graph Neural Networks using Localization

### v3

@@ 9.3 Implementation Details
We experimented with both InsideOutGNN and InsigGNN models to predict the count of structures such as triangles, 3stars, 2stars, and Chordal cycles. For each of these structures, the true local counts are computed using various methods. For example, if our task is to count the number of triangles in a particular subgraph, for the root node of the subgraph, we label its local count to be the number of edges present in a 1-hop neighbouring subgraph without the root node. Using these local counts, from Equation Lemma 5, we assign the global counts to be $\frac{1}{3}\sum\limits_{u\in V(G_{v}^{1})}\prescript{x\mkern-0.5mu}{}{C}_{2}$, where $x$ is the local count corresponding to the root vertex $v.$ Here, $3$ is the orbit of a node in a triangle. Similarly, we assign the local counts and the global counts for different substructures.
When counting larger substructures such as $K_{4}$ and $C_{4},$, we use the fragmentation technique with the help of the InSignGNN model. For calculating $K_{4}$, we construct a set of subgraphs. For every node in the graph, a 1-hop neighbouring subgraph is generated, from which we remove the root node. In this subgraph, we count the number of triangles and process accordingly. Therefore, we use the model already trained for counting triangles in order to count $K_{4}.$ During the inference, we apply a rounding function, as the counts are integer numbers.
@@ 9.4 Hyperparameters
We use two GIN Convolutional layers for the Internal GNN part in both proposed models. For the models we have considered, as there can be paths of lengths more than 1 in the subgraph, $2$ GINConv layers are sufficient to capture the information well. In the InsideOut model, we use $2$ GINConv layers in the external part of the model as well.
We use a learning rate of $41e-$ 4 and a batch size of $1$ We also experimented with different hidden dimensions for the node embeddings and obtained the best results when we used $4$ as a hidden dimension size. The ablation studies have been mentioned in Section 10.2. The experiments were conducted using an NVIDIA A100 40GB GPU.
10 Experimental Observations
10.1 Results
Without Fragmentation Fragmentation
Models Triangle 3-star 2-star Chordal $K_{4}$ $C_{4}$ Tailed triangle
ID-GNN 6.00E-04 NA NA 4.54E-02 2.60E-03 2.20E-03 1.05E-01
@@ 10.1 Results
I2-GNN 3.00E-04 NA NA 1.00E-03 3.00E-04 1.60E-03 1.10E-03
InsideOut 0.00E+00 0.00E+00 0.00E+00 3.82E-03
InSig 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Table 4: MAE for the subgraph count of different patterns. Some results, such as $2-$star, $3-$star, are not conducted by the given baselines; therefore, it is mentioned NA. As mentioned in Section Section 10.1, InSig Model has a lesser number of parameters while giving the same performance as InsideOut Model; therefore, we used only InSig Model for fragmentation based counting.
We conducted our experiments***The dataset and the code are present in the anonymous repository[Link]. for counting the following patterns: 1. triangles, 2. $3$stars, 3. $2$stars, 4. chordal cycles, 5. $K4,$ 6. $C4$ and 7. tailed triangles. The dataset used for the experiments is a random graphs dataset prepared in Chen et al. (2020). We report our experiments’ Mean Absolute Error (MAE) in Table 4. We compare our results with those reported in ID-GNN You et al. (2021), NGNN Zhang and Li (2021), GNNAK+ Zhao et al. (2022), PPGN Maron et al. (2019) and I2-GNN Huang et al. (2023). We can observe that our approach of predicting the local counts first and then predicting the global counts, achieves zero error for almost all of the tests. Although we experimented with two models, we realise that InSig model needs much fewer parameters to learn yet gives the same results. For the structures where we used fragmentation, we apply the models for counting triangles and 2stars in order to count $K_{4}$ and $C_{4}$, respectively. Also, we use the InSig model to predict the degree of a node and the number of edges in the induced subgraph without the root node. We call them the Local Node and Local Edge models, respectively. Using the Local Node and Local Edge models, we count the number of Tailed Triangles.
For all the patterns, we observed that the model gets to zero error after only 2 to 3 epochs.
@@
Inference Time
Memory Usage
@@
InSig 190 20 1.2
Table 5: The table shows the comparison of the number of parameters required by the baselines and InSig Model. The values shown here correspond to the triangle counting task with the hidden dimension set as $4$. It can be observed that our approach requires a considerably smaller number of parameters and beats all the baselines. Our inference time comprises model inference as well as the graph preprocessing time, where it creates the set of subgraphs corresponding to each node in the graph.
10.2 Ablation Study
We varied the hidden dimension of the GINConv layer by taking the hidden dimensions from the following set $\{2,4,8,16,32,64\}$. The loss corresponding to different hidden dimensions is plotted in Figure 5. The experiment was conducted using the InSig Model for counting triangles. We observe that our model was able to compute exact counts with just 4 hidden dimensions.
Figure 5: Ablation Study - Validation Loss VS Hidden Dimension
We also varied the number of GINConv Layers. The observations are present in Table 6. The task for this ablation study was counting triangles. As per the previous study, we set the hidden dimension as $4$. We observe that we two GINConv Layers we get the best result. It should also be noted that with other sets of layers, the loss is better than the baselines.
Number
of
Layers
Number of
Internal
Model
Parameters
Number of
External
Model
Parameters
Total
Parameters
Validation
Loss
Test
Loss
1 53 97 150 1.41E-04 4.24E-04
2 93 97 190 0.00E+00 0.00E+00
3 133 97 230 2.35E-05 9.42E-05
Table 6: This table shows the ablation study with the number of GINConv Layers in the proposed architecture.
The fragmentation technique helps us to use the previously learned models, such as the model to count triangles to count the number of $K_{4}$ in a graph. Although the loss corresponding to that experiment is zero, as mentioned in Table 4, we did another study where we used a different dataset to train previous models of fragmentation. For example, we trained the InSig model to count triangle and 2stars on $5000$ Erdos-Renyi random graphs Chen et al. (2020). Using the respective trained models, we predicted $K_{4}$ and $C_{4}$ counts in the random graphs dataset prepared in Chen et al. (2020). The reverse study was also conducted. In both cases, the loss was zero. This shows that our proposed architecture is not dependent on the statistics of the dataset. The fragmentation shows that a trained model to count the number of smaller patterns on any dataset can be used to count the larger patterns.
@@ 11 Conclusion
In this work, we progressed toward a more precise characterization of the localized versions of WL algorithms. We showed how the Local-$k-$WL lies between the global $k$ and $k+1-$WL in terms of expressiveness. We also developed strategies to make the Local-$k-$WL algorithm more efficient by introducing techniques such as layered WL, recursive WL, and pattern fragmentation. The hope is that such generalizations of the WL algorithm will lead to a finer subdivision of the WL hierarchy as well as more efficient and expressive graph neural networks.

### v5

@@ 9.1 Implementation Details
We refer to our proposed model as InSig. In the experiments, we predict the counts of the following substructures occurring as subgraphs : triangles, $3$-Star, 22-Star, chordal $C_{4}$, $K_{4}$, $C_{4}$, and tailed triangles.
Task
Total pattern
count
Zero Count
graphs
Standard Deviation
of Count
Average number
of Nodes
Average number
of Edges
Number
of graphs
Triangle $25209$ $195$ $3.072$ $18.7976$ $62.678$ $5000$
2-Star $429463$ 0 $18.015$
3-Star $309525$ 0 $17.777$
Chordal $19088$ $1786$ $4.742$
K4 $643$ $4447$ $0.387$
C4 $53002$ $16$ $6.938$
Tailed Triangle $177968$ $195$ $25.943$
Table 3: Dataset statistics. The total number of graphs in the dataset is $5000.$ We used $4000$ graphs for training, $500$ for validation and $500$ for testing.
When counting larger substructures like $K_{4}$, $C_{4}$, and Tailed Triangles, we use the fragmentation technique with the help of a model learned to predict triangles, 3-Star, 2-Star, and Chordal $C_{4}$. For counting $K_{4}$, the task of Pattern Counting Component is to learn which edges to prune. Later Local Count Learning component counts the number of substructures, which in the case of $K_{4}$, is triangles. Once the number of triangles is predicted using models learned to count triangles, the normalization factor is learned to output the global count. In other structures like tailed triangles, we have two patterns to learn: the nodes in the $1$-hop neighborhood and the edges between them. During the inference phase, we use a rounding function, as the counts are integer numbers.
@@ 9.2 Hyperparameters
We use two GIN Convolutional layers for the Pattern Learning Local Count Learning component. We consider Linear transformations as readout layers in the previously mentioned components. Since we need to classify whether an edge should be present or not in a subgraph for counting a subpattern, we use Binary Cross-Entropy (BCE) loss for the pattern learning component. For the local counting and global counting components, we use Mean Absolute Error (MAE). For the models we have considered, as there can be paths of lengths more than 1 in the subgraph, $2$ GINConv layers are sufficient to capture the information well.
We use a learning rate of 1e41e-4 and a batch size of $1.$ We also experimented with different hidden dimensions for the node embeddings and obtained the best results when we used $4$ as a hidden dimension size. The experiments were conducted using an NVIDIA A100 40GB GPU. The source code of the implementation is available at this Github Link.
9.3 Experimental Results
Models Without Fragmentation Fragmentation
Triangle 3-Star 2-Star Chorcal C4 $K_{4}$ $C_{4}$ Tailed Triangle
ID-GNN 6.00E-04 NA NA 4.52E-02 2.60E-03 2.20E-03 1.05E-01
@@ 9.3 Experimental Results
I2-GNN 4.00E-04 NA NA 1.00E-03 3.00E-04 1.60E-03 1.10E-03
InSig 0.00E-$0$ 0.00E-$0$ 0.00E-$0$ 0.00E-$0$ 0.00E-$0$ 0.00E-$0$ 0.00E-$0$
Table 4: MAE for the subgraph count of different patterns. Some results, such as 22-star, 33-star, are not conducted by the given baselines; therefore, it is mentioned NA.
The dataset used for the experiments is a random graphs dataset prepared in (chen2020can). In Table 3, we report the dataset statistics, specifically the counts of the various patterns used for our experiments, number of graphs where these patterns do not appear, and so on. We report our experiments’ Mean Absolute Error (MAE) in Table 4. We compare our results with those reported in ID-GNN (https://doi.org/10.48550/arxiv.2101.10320), NGNN (zhang2021nested), GNNAK+ (zhao2022from), PPGN (https://doi.org/10.48550/arxiv.1905.11136) and I2-GNN (huang2023boosting). We can observe that our approach of predicting substructures, the local counts, and then predicting the global counts, achieves zero error for all of the tests.
For all the patterns, we observed that the model gets to zero error after only 2 to 3 epochs. From Table 5, it can be observed that our approach requires a considerably smaller number of parameters and beats all the baselines. Our inference time comprises model inference as well as the graph preprocessing time, where it creates the set of subgraphs corresponding to each node in the graph.
@@
Inference
Time (ms)
Memory
Usage (GB)
@@
InSig 268 20 1.2
Table 5: The table shows the comparison of the number of parameters required by the baselines and InSig Model. The values shown here correspond to the triangle counting task with the hidden dimension set as $4$.
(a) Figure shows an example of $G^{1}_{0}$ for a random graph with $10$ nodes
(b) t-SNE plot of the updated node embeddings from the Pattern learning component
Figure 4: Figure showing the input and updated node embedding from the pattern learning component
In Figure 4, we show an example of a graph sent as input to the pattern learning component and the updated node embeddings outputted from the model. In Figure 4(a), we see that the root node is node 0 and there are directed edges from the root node to its neighboring nodes. Figure 4(b) shows the updated node embeddings of each node in the graph. We can observe that $8,2,4,6$ nodes can be separated using some separator from the rest of the nodes. This indicates that the model is able to learn the updated node embedding such that we can distinguish nodes which has edges between them.
@@ 10 Conclusion
Subgraph counting is a fundamental combinatorial problem that arises in the study of graphs and graph structured problems as well as data. Exactly counting the number of subgraphs in a graph is also a computationally hard problem. In this paper, we present a learnable algorithm that is able to compute exact counts of a number of commonly occurring patterns. The proposed fragmentation method has proven to be beneficial for the task of counting subgraphs. Additionally, since fragmentation results in smaller subgraphs, the parameter requirements of the proposed GNN-based architecture is orders of magnitude less than previous methods. As future work, we plan to analyse the fragmentation algorithm from a theoretical perspective. We also plan to investigate the effectiveness of this technique for increasing the expressiveness of GNNs for downstream tasks. Along with that, the current approach is limited to patterns that can be computed using $1$-hop or 22-hop subgraphs, hence we can study our algorithm for bigger patterns.